Name a US government agency that publishes data of CPI and unemployment and provide a URL...

  1. Name a US government agency that publishes data of CPI and unemployment and provide a URL link for each of the two variables that leads you to the data.
  2. Calculate the inflation rate between year a) 1999 and 2000, b) 2000 and 2001 given the following CPI:

CPI (1999)= 89, CPI (2000)=100, CPI (2001)=107

  1. Suppose the nominal output of the year 1999 was $3200 and the real output of the year 1999 was $3600, calculate the GDP deflator of year 1999.
  2. Suppose the CPI in Maryland is 110 and the CPI in Texas 1999 is 100 and you are given that the nominal income in Texas is $5000 while the nominal income in Maryland is $5100. Suppose you are given a job offer in each state. Which offer should you take if you are indifferent about the different environment in each state? Explain your answer thoroughly and include necessarily calculations or numbers.
  3. Suppose the CPI of the US in 1990 was 120 and the CPI in 2003 was 200. For how much could a car that cost $20,000 in 2003 have been bought in 1990? Explain your answer thoroughly and include necessarily calculations or numbers.
  4. If the real interest rate is 5% and inflation rate is 3 %, what is the nominal interest rate?
Answer #1

Ans A)

Inflation Rate in between 1999 and 2000

=CPI(2000)-CPI(1999)/CPI(1999)=100-89/100=11%

INflation rate in between 2000 and 2001

=CPI(2001)-CPI(2000)/CPI(2000)=107-100/100=7%

Ans B)

GDP Deflator=Nominal GDP/Real GDP=3200/3600=8/9=0.89=89

Ans C)

Real Income=Nominal Income/Price Index

Real Income(Texas)=5000/1=$5000

Real Income( MAryland)=5100/1.1=$4636

Real Income in Texas is higher than Real Income in Maryland therefore he should choose to work in Texas

Ans D)

120(1+r)^13=200

r=(200/120)^(1/13)-1=4.00%

X(1.04)^13=20000

X=20000/(1.04)^13\=12011.48

Ans E)

Nominal Rate=Real Rate+Inflation Rate=5%+3%=8%